# Attributed-graphs kernel implementation using local detuning of neutral-atoms Rydberg Hamiltonian

### v1

@@ Appendix C Expressiveness of GDQC
where $l\leq l_{\mathrm{max}}$. The base case is $l=l_{\mathrm{max}}$ which is exactly our Eq( 43). Now assume that (44) holds, when colors at round $l$ are computed, vertex $i$ in $\mathcal{G}_{1}$ hashes $R_{\mathcal{G}_{1}}^{(l)}\left(i\right)$. By the induction hypothesis we have $R_{\mathcal{G}_{1}}^{(l-1)}\left(i\right)=R_{\mathcal{G}_{2}}^{(l-1)}\left(\Pi\left(i\right)\right)$ which means that each ordered pair in $R_{\mathcal{G}_{1}}^{(l-1)}\left(i\right)$ occurs with the same frequency in $R_{\mathcal{G}_{2}}^{(l-1)}\left(\Pi\left(i\right)\right)$. This allows us to align each tuple with distance $D_{i,j}^{\mathcal{G}_{1}}$ with the corresponding tuple $D_{\Pi\left(i\right),\Pi\left(j\right)}^{\mathcal{G}_{2}}$, as the distances don’t change from one round to the other, ensuring that $\chi_{\mathcal{G}_{1}}^{(l-1)}\left(j\right)=\chi_{\mathcal{G}_{2}}^{(l-1)}\left(\Pi\left(j\right)\right)$ valid for $\forall j\in\mathcal{V}_{1}$.
@@
From here, the proof of proposition 3 is straightforward. If two graphs are undistinguishable by the GD-WL test with $q_{t}\left(i\right)$ as initial coloring, then even when the number of bins is large enough (with maximal expressiveness attained when $N_{\mathrm{bins}}^{D}>\left[\min\left(D^{\mathcal{G}}_{i,j}-D^{\mathcal{G}}_{i^{\prime},j^{\prime}}\right)\right]^{-1}$ and $N_{\mathrm{bins}}^{C}>\left[\min\left(C_{i,j}-C_{i^{\prime},j^{\prime}}\right)\right]^{-1}$), according to lemma 5 their corresponding tensors $M^{\mathcal{G}_{1}}$ and $M^{\mathcal{G}_{2}}$ are one permutation away one from the other. Considering that hash is an injective function, this implies that, at a fixed graph distance value, the multi-sets obtained from the corresponding correlation matrix rows are also the same, and thus the number of times a given value for the correlation appears (while its corresponding nodes are at the same fixed distance) is also the same. Given the definition of the GDQC kernel, the corresponding feature vectors for such a pair of graphs are identical, and their kernel value is equal to $1$, which completes the proof. ∎
Appendix D Kernel pooling statement
In this appendix, we demonstrate the claim of the proposition 4: For a set $S=\{\kappa_{1},\kappa_{2},...,\kappa_{N}\}$ of graph kernels, let $\kappa_{A}$ and $\kappa_{B}$ be respectively the kernels obtained from pooling two subsets $S_{A}$ and $S_{B}$ of $S$. If $S_{A}\subset S_{B}$ then $\kappa_{A}$ is at least as expressive as $\kappa_{B}$.

### v2

@@ Appendix E Expressiveness of $\mathrm{GDQC}$
where $l\leq l_{\mathrm{max}}$. The base case is $l=l_{\mathrm{max}}$ which is exactly our (73). Now assume that (74) holds, when colors at round $l$ are computed, vertex $i$ in $\mathcal{G}_{1}$ hashes $R_{\mathcal{G}_{1}}^{(l)}\left(i\right)$. By the induction hypothesis we have $R_{\mathcal{G}_{1}}^{(l-1)}\left(i\right)=R_{\mathcal{G}_{2}}^{(l-1)}\left(\Pi\left(i\right)\right)$ which means that each ordered pair in $R_{\mathcal{G}_{1}}^{(l-1)}\left(i\right)$ occurs with the same frequency in $R_{\mathcal{G}_{2}}^{(l-1)}\left(\Pi\left(i\right)\right)$. This allows us to align each tuple with distance $D_{i,j}^{\mathcal{G}_{1}}$ with the corresponding tuple $D_{\Pi\left(i\right),\Pi\left(j\right)}^{\mathcal{G}_{2}}$, as the distances don’t change from one round to the other, ensuring that $\chi_{\mathcal{G}_{1}}^{(l-1)}\left(j\right)=\chi_{\mathcal{G}_{2}}^{(l-1)}\left(\Pi\left(j\right)\right)$ valid for $\forall j\in\mathcal{V}_{1}$.
@@
From here, the proof of proposition 3 is straightforward. If two graphs are undistinguishable by the GD-WL test with $q_{t}\left(i\right)$ as initial coloring, then even when the number of bins is large enough (with maximal expressiveness attained when $N_{\mathrm{bins}}^{D}>\left[\min\left(D^{\mathcal{G}}_{i,j}-D^{\mathcal{G}}_{i^{\prime},j^{\prime}}\right)\right]^{-1}$ and $N_{\mathrm{bins}}^{C}>\left[\min\left(C_{i,j}-C_{i^{\prime},j^{\prime}}\right)\right]^{-1}$), according to lemma 5 their corresponding tensors $M^{\mathcal{G}_{1}}$ and $M^{\mathcal{G}_{2}}$ are one permutation away one from the other. Considering that hash is an injective function, this implies that, at a fixed graph distance value, the multi-sets obtained from the corresponding correlation matrix rows are also the same, and thus the number of times a given value for the correlation appears (while its corresponding nodes are at the same fixed distance) is also the same. Given the definition of the $\mathrm{GDQC}$ kernel, the corresponding feature vectors for such a pair of graphs are identical, and their kernel value is equal to $1$, which completes the proof. ∎
Appendix F Kernel pooling statement
In this appendix, we demonstrate the claim of the proposition 4: Let $S=\{\kappa_{1},\dots,\kappa_{N}\}$ be a set of graph kernels, and let $\kappa_{A}$ and $\kappa_{B}$ be kernels obtained by pooling two subsets $S_{A}$ and $S_{B}$ of $S$ using (19) or (20) as pooling rules. If $S_{A}\subset S_{B}$, then $\kappa_{B}$ is at least as expressive as $\kappa_{A}$.